Cyberdyne Systems Corporation uses a standard-costing system. The production budget for the fiscal year (FY) ended 30 November 2018 was based on 200,000 units of output. Each unit requires two standard hours of manufacturing labor for completion. Total overhead was budgeted at $900,000 for the year, and the budgeted fixed overhead rate was $3 per unit of output. Both fixed and variable overhead are allocated to the product on the basis of direct manufacturing labor-hours. The actual data for the fiscal year ended 30 November 2018 are as follows: actual production in units, 198,000; actual direct manufacturing labor-hours, 440,000; actual variable overhead, $352,000; actual fixed overhead, $575,000.

The Cyberdyne fixed overhead spending variance for FY2018 is closest to:

a.

$19,000 F

b.

$19,000 U

c.

$25,000 F

d.

$25,000 U

The Cyberdyne output level variance for FY2018 is closest to:

a.

$25,000 F

b.

$55,000 U

c.

$19,000 F

d.

$6,000 U

The Cyberdyne variable overhead spending variance for FY2018 is closest to:

a.

$20,000 U

b.

$20,000 F

c.

$22,000 U

d.

$22,000 F

The Cyberdyne variable overhead efficiency variance for FY2018 is closest to:

a.

$33,000 U

b.

$35,520 F

c.

$66,000 U

d.

$35,200 U
Answer #1

Solution:

working notes:

First calculate the budgeted fixed and variable overhead

budgeted fixed overhead = $3 per unit of output

so budgeted fixed overhead = 200,000 units \times $3 per unit

= $600,000

Total budgeted overhead = $900,000

Budgeted Variable Overhead = Total budgeted overhead - Budgeted Fixed Overhead

   = $900,000 - $600,000

= $300,000

now we calculate standard hours for actual output

Standard Hours for Actual Output = 198,000 units \times 2 hours per unit

=$396,000

Now we calculate Budgeted overhead rate

Budgeted overhead rate =Budgeted Overhead \div Budgeted hours

= $900,000 \div *400,000 hours

= $2.25 per hour

*200,000 units \times 2 hours per unit = 400,000 hours

1.  Cyberdyne fixed overhead spending variance for financial year 2018 is closest to

Fixed overhead spending variance = Budgeted fixed overhead - Actual fixed Overhead

  =$600,000 - $575,000

= $25,000(F)

so (C) option is correct

2. Cyberdyne Output level Variance for FY 2018

Output level Variance = (Actual unit Produced - Budgeted Unit Produced) \times Budgeted Overhead Rate

= (198,000 units - 200,000 units) \times $2.25

   = $4,500(U)

so (D) option is correct

According to options The Closest answer is $6,000(U) but the exact answer is $4,500(U)

3. Cyberdyne Variable Overhead Spending variance for FY 2018

Variable Overhead Spending Variance = (Standard variable absorption rate - Actual Variable Rate) \times Actual hours

= (0.75a - 0.80b) \times 440,000 hours

= -0.05 \times 440,000 hours

= $22,000 (U)

a. $300,000 \div 400,000 hours = $0.75

b. $352,000 \div 440,000 hours = $0.80

Option (C) is correct

4. cyberdyne Variable Overhead Effeciency Variance

Variable Overhead Effeciency Variance = (Standard Hours for Actual Output - Actual Hours ) \times standard variable overhead rate

= (396,000 units -440,000 units) \times $0.75

= - 44,000 units \times $0.75

= $33,000(U)

so (A) Option is Correct